OpenStudy (helpblahblahblah):

ILL GIVE MEDALS: Please explain Wildlife officials had to rescue and remove 735 Florida panthers in the past 12 months. It is estimated that there are 50,000 Florida panthers in existence. WHAT PERCENT OF 50,000 IS 735? A. 1.47% B. 7.35% C. 36.75% D. 68.02%

OpenStudy (mertsj):

Do the translation: what percent: x/100 Of: times Is: =

OpenStudy (studygurl14):

What percent of 50,000 is 175? of = multiply is = equal to WP * 50,000 = 175 Solve for WP

OpenStudy (helpblahblahblah):

im confused?

OpenStudy (studygurl14):

Then times by 100 and add %

OpenStudy (studygurl14):

WP x 50,000 = 175 WP = 175 ÷ 50,000 WP = 0.0035 0.0035 x 100 = 0.35, meaning 0.35%

OpenStudy (helpblahblahblah):

okay thank you

OpenStudy (studygurl14):

you're welcome :)